OpenStudy (anonymous):

Hyperplanes? Find a basis for the hyperplane vector A perp? A=(1,2,-3)

OpenStudy (anonymous):

The answer is {(-2,1,0),(3,0,1)} if it helps.. I really don't get how they got this

OpenStudy (amistre64):

doesnt ring a bell, srry

OpenStudy (anonymous):

A is the cross of the answers....

OpenStudy (anonymous):

U would go (a-b)X(a-c) to get a normal.

OpenStudy (zarkon):

that is an answer...there are infinity many answers to your question.

OpenStudy (zarkon):

(2,-1,0),(0,3,2) is a second answer

OpenStudy (anonymous):

Let \(v_1 = A\) then select two other vectors \(v_1, v_2\) such that \(\{v_1, v_2, v_3\}\) is linearly independent.Then perform the Gram-Schmidt orthogonalization process.

OpenStudy (zarkon):

total overkill

OpenStudy (zarkon):

just pick any two vectors that are linear independent and orthogonal to the vector given

OpenStudy (zarkon):

for the vector given (1,2,-3) make a vector that has one coordinate zero (x,y,0) now switch the 1 and 2 and make one of them negative (-2,1,0) you just found an orthogonal vector now create another vector with the zero in a different location (a,0,b) switch 1 and -3 and change the sign of on of them (3,0,1) done!

OpenStudy (anonymous):

Can I make the first one 0 then switch 2 and -3?

OpenStudy (anonymous):

are there other possiblilities is what I'm trying to say, and is that how you go about for a hyperplane

OpenStudy (zarkon):

there are infinity many possibilities...I think this is the easiest way to do it in this case

OpenStudy (anonymous):

I am more used to it the other way about. U have abc in the plane (in R3) and u get p = b-a and q = c-b then n =pXq then if u dot that with a (in xyz) u get equation for a plane.

OpenStudy (zarkon):

"Can I make the first one 0 then switch 2 and -3?" yes...see my first answer

OpenStudy (anonymous):

Oh okay thank you! Estudier I get what you're saying but what's with the X in pXq

OpenStudy (zarkon):

the X is for the cross product

OpenStudy (anonymous):

Oh okay so once I get the equation for a plane how would I find the basis?

OpenStudy (anonymous):

Sorry for the questions just studying for my final and this is like the one concept I don't quite understand

OpenStudy (zarkon):

pick any two linearly independent vectors in you plane...they will form a basis

OpenStudy (anonymous):

I like perpendicular ones...

OpenStudy (zarkon):

those are nice...but you don't need them unless they specifically ask for vectors that are orthogonal

OpenStudy (anonymous):

Okay so if there's more like 3 values, like a=(2,-1,4,1), I can do the same thing but I'd have three vectors?

OpenStudy (anonymous):

more than 3*

OpenStudy (anonymous):

just make each part 0 and switch the values but would one still need to be negative

OpenStudy (anonymous):

cross is a problem in 4d

OpenStudy (anonymous):

okay I'm lost again haha so how would do it with four

OpenStudy (zarkon):

{(1,2,0,0),(0,4,1,0),(0,0,-1,4)} would work

OpenStudy (zarkon):

they are linearly independent and they are all orthogonal to the vector you gave above

OpenStudy (anonymous):

This sort of thing is why I switched to Geometric Algebra...

OpenStudy (anonymous):

the answer says (0,1,0,1) (0,0,1,-4),(0,4,1,0)

OpenStudy (anonymous):

how are you getting this zarkon?

OpenStudy (zarkon):

that is one answer...there are infinity many...as I stated before

OpenStudy (zarkon):

the same way your book probably did it...make to entries 0 and switch the other two and change the sign on one of them

OpenStudy (zarkon):

make two entries...

OpenStudy (anonymous):

Then why does some of them have two zeroes.. My book doesn't explain hyperplanes like at all

OpenStudy (anonymous):

Oh snap I reread your thing

OpenStudy (zarkon):

both your answer and mine have exactly two zeros in each vector.

OpenStudy (anonymous):

Okay if it was in 5 d, I'd make 3 entries zeroes?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Oh okay I will try this thanks for clearing it, and thanks to everyone else who answered!

OpenStudy (anonymous):

Eventually you will see the normal equation description of a hyperplane which makes life a lot simpler...

OpenStudy (zarkon):

the answer you gave does not work

OpenStudy (zarkon):

(0,1,0,1) (0,0,1,-4),(0,4,1,0) is not a basis

OpenStudy (zarkon):

4(0,1,0,1)+(0,0,1,-4)=(0,4,1,0)

OpenStudy (anonymous):

? That's the answer in my book {(0,1,0,1) (0,0,1,-4),(0,4,1,0)}

OpenStudy (zarkon):

it's wrong

OpenStudy (zarkon):

those vectors are dependent and therefore do not form a basis

OpenStudy (anonymous):

Okay I understand that a basis needs to be independent and span all of R^n. So is that what a hyperplane is? Are we just taking the vector and find other vectors that can form a basis?

OpenStudy (anonymous):

nvm I understand it now. Thanks for your help Zarkon